OpenStudy (laurenarg):

Enter the equation of the circle described below. Center (0, 4), radius sqrt 3

OpenStudy (peachpi):

The equation of a circle is \[(x-h)^2+(y-k)^2=r^2\]where (h, k) is the center and r is the radius. Plug in your numbers

OpenStudy (laurenarg):

What about the sqrt 3 thing... I don't understand this at all

OpenStudy (peachpi):

\(\sqrt{3}\) is the radius, so plug that in for r

OpenStudy (peachpi):

Look at the center provided in the question. What are h and k?

OpenStudy (laurenarg):

(x)2 + (y-4)2

OpenStudy (peachpi):

make sure you use ^ for exponents. (x)^2 + (y-4)^2 is correct for the left side. Now you need the right side. Plug in √3 for r. What is \[\left( \sqrt{3} \right)^2\]

OpenStudy (laurenarg):

3??

OpenStudy (peachpi):

right, so the equation of the circle is\[x^2+(y-4)^2=3\]

OpenStudy (laurenarg):

Thank you so much!

OpenStudy (peachpi):

you're welcome